AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 235     Entry time: Tue Feb 22 12:01:46 2022
Author: RSS, OF, ML, CB 
Subject: Beam OFF, Target ON Run 54  

01:00

Beam optimisation since 09:25.

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

Beam optimization will be done until 3 pm and will be then give for the experiment.

Attachment 1: 12-ex-2022-02-22_13-00-14.png  153 kB  | Hide | Hide all
12-ex-2022-02-22_13-00-14.png
Attachment 2: 12-ey-2022-02-22_13-00-40.png  154 kB  | Hide | Hide all
12-ey-2022-02-22_13-00-40.png
Attachment 3: 12-xy-2022-02-22_13-01-10.png  195 kB  | Hide | Hide all
12-xy-2022-02-22_13-01-10.png
Attachment 4: 12-adc-2022-02-22_12-59-08.png  102 kB  | Hide | Hide all
12-adc-2022-02-22_12-59-08.png
Attachment 5: 12-good-2022-02-22_12-58-48.png  104 kB  | Hide | Hide all
12-good-2022-02-22_12-58-48.png
Attachment 6: 12-temp-2022-02-22_12-57-28.png  113 kB  | Hide | Hide all
12-temp-2022-02-22_12-57-28.png
Attachment 7: 12-leakage-2022-02-22_12-59-32.png  21 kB  | Hide | Hide all
12-leakage-2022-02-22_12-59-32.png
Attachment 8: 12-spectrum-2022-02-22_12-58-02.png  177 kB  | Hide | Hide all
12-spectrum-2022-02-22_12-58-02.png
Attachment 9: 12-correlation-2022-02-22_12-58-25.png  102 kB  | Hide | Hide all
12-correlation-2022-02-22_12-58-25.png
ELOG V3.1.4-unknown